# Measurability in the Fundamental Theorem of Statistical Learning

### v2

@@ 3 Fundamental Theorem of Statistical Learning
The aim of this section is to establish the Fundamental Theorem of Statistical Learning. This result, due to Blumer, Ehrenfeucht, Haussler and Warmuth [5, Theorem 2.1 (i)], beautifully relates the notions VC dimension and PAC learnability. More precisely, it states that, under certain broadly applicable measurability conditions, a hypothesis space is PAC learnable if and only if its VC dimension is finite. In the proofs presented in this article, we follow modern expositions such as Anthony and Bartlett [2], Shalev-Schwartz and Ben-David [27] and Vidyasagar [36]. We place particular emphasis on the required measurability assumptions, both in the statements and in the proofs of the results.
@@
$\underline{z}$ $\mapsto U(\underline{z}):=\sup\limits_{h\in\mathcal{H}}\big{|}\operatorname{er}_{\mathbb{D}}(h)-\hat{\operatorname{er}}_{\underline{z}}(h)\big{|}$
@@
The proof of the Fundamental Theorem of Statistical Learning presented below, specifically the proof of Theorem 3.6, relies on computing and bounding the expected values of the maps $U$ and $V$. Therefore, the measurability of these maps is necessary for our arguments to apply. Sufficient conditions for hypothesis spaces to be well-defined can be found in Appendix A.2.
@@
At this point, we emphasize that in the literature there are several extended versions of this theorem. These entail further equivalences and quantitative aspects such as sample complexity bounds (see e.g. [2, Theorem 5.5], [5, Theorem 2.1], [27, Theorem 6.7 and Theorem 6.8]). The main aim for the remainder of this section is to establish the Fundamental Theorem of Statistical Learning as stated above by virtue of the following more extended version:
@@
In the following we establish Theorem 3.3 by providing self-contained proofs of the implications (1$\Rightarrow$(2), (2$\Rightarrow$(3), (3$\Rightarrow$(4) and (4$\Rightarrow$(1). Our arguments for the implications (3$\Rightarrow$(4) and (4$\Rightarrow$(1) also apply to hypothesis spaces that are not well-behaved. However, for the uniform convergence property to be well-defined, the map $U$ from Definition 2.4 has to be measurable. Furthermore, our proof of implication (1$\Rightarrow$(2) is based on the assumptions that both the maps $U$ and $V$ are random variables, as we compute and bound their expected values, respectively. For our proof of implication (4$\Rightarrow$(1), the set $\mathcal{D}$ needs to contain all discrete uniform distributions. This assumption is not necessary for the arguments we use to establish the other implications.
(1$\Rightarrow$(2)
@@
Let $\mathcal{X}$ be a non-empty set, let $\Sigma_{\mathcal{Z}}$ be a $\sigma$–algebra on $\mathcal{Z}=\mathcal{X}\times\{0,1\}$ and let $h\in\{0,1\}^{\mathcal{X}}$ with $\Gamma(h)\in\Sigma_{\mathcal{Z}}$. Then the hypothesis space =$\mathcal{H}=\{h\}$ has the uniform convergence property with respect to any set $\mathcal{D}$ of distributions on $(\mathcal{Z},\Sigma_{\mathcal{Z}})$.
@@
$\mathbb{D}^{m}\left(\left\{\underline{z}\in\mathcal{Z}^{m}\left|\ \big{|}\operatorname{er}_{\mathbb{D}}(h)-\hat{\operatorname{er}}_{\underline{z}}(h)\big{|}>\varepsilon\right.\right\}\right)\leq\delta.$
@@
$\big{|}\operatorname{er}_{\mathbb{D}}(h)-\hat{\operatorname{er}}_{\underline{z}}(h)\big{|}=\frac{1}{m}\big{|}Y_{1}(\underline{z})+\dots+Y_{m}(\underline{z})\big{|},$
@@
is a random variable [^8] The map $Y_{i}$ is a translation of the map $\theta_{i}\colon\mathcal{Z}^{m}\to\{0,1\},\ \underline{z}\mapsto\ell(h,z_{i})$, which is $\Sigma_{\mathcal{Z}}^{m}$–measurable, as $\Gamma(h)\in\Sigma_{\mathcal{Z}}$ implies $\theta_{i}^{-1}(\{0\})=\mathcal{Z}\times\dots\times\mathcal{Z}\times\underbrace{\Gamma(h)}_{i\text{--th position}}\times\mathcal{Z}\times\dots\times\mathcal{Z}\in\Sigma_{\mathcal{Z}}^{m}.$ with $\mathbb{E}_{\underline{z}\sim\mathbb{D}^{m}}[Y_{i}(\underline{z})]=0$ for any $i\in[m]$. Straightforward computations show that the random variables $Y_{1},\dots,Y_{m}$ are independent. Thus, we can apply Hoeffding’s Inequality (cf. Pollard [25, Appendix B, Corollary 3]) to obtain
$\mathbb{D}^{m}\left(\left\{\underline{z}\in\mathcal{Z}^{m}\left|\ \big{|}\operatorname{er}_{\mathbb{D}}(h)-\hat{\operatorname{er}}_{\underline{z}}(h)\big{|}>\varepsilon\right.\right\}\right)$ $\leq 2\exp(-m\varepsilon^{2}/2)$
@@
We next prove a result that shows how the growth function $\pi_{\mathcal{H}}$ can be used to bound the probability involved in the definition of the uniform convergence property. The proof requires the hypothesis space to be well-behaved, since its core lies in bounding the expected values of the random variables $U$ and $V$. The original proof of Vapnik and Chervonenkis is also based on the assumption that these maps are measurable (see [34, pages 265 and 268]). The proof presented here is based on the one of [27, Theorem 6.11]. We present here a modified and adapted version of [27, proof of Theorem 6.11], highlighting explicitly where the required measurability assumptions come into play.
@@
$\mathbb{D}^{m}\left(\left\{\underline{z}\in\mathcal{Z}^{m}\left|\ \sup\limits_{h\in\mathcal{H}}\big{|}\operatorname{er}_{\mathbb{D}}(h)-\hat{\operatorname{er}}_{\underline{z}}(h)\big{|}\leq\varepsilon_{0}(m,\delta)\right.\right\}\right)\geq 1-\delta,$
@@
$\underline{z}$ $\mapsto U(\underline{z}):=\sup\limits_{h\in\mathcal{H}}\big{|}\operatorname{er}_{\mathbb{D}}(h)-\hat{\operatorname{er}}_{\underline{z}}(h)\big{|}.$
We have to show that $\mathbb{D}^{m}(\{\underline{z}\in\mathcal{Z}^{m}\mid U(\underline{z})\leq\varepsilon_{0}(m,\delta)\})\geq 1-\delta$. Since the random variable $U$ is non-negative, by Markov’s Inequality (see [27, Appendix B.1]) it suffices to establish that
@@
writing $\underline{z}=(z_{1},\dots,z_{m})$ and $\underline{z}^{\prime}=(z^{\prime}_{1},\dots,z^{\prime}_{m})$. The expected value in (3.2) is determined based on a choice of multi-sample $(\underline{z},\underline{z}^{\prime})\in\mathcal{Z}^{2m}$ whose components are independently distributed according to $\mathbb{D}$, i.e. $(\underline{z},\underline{z}^{\prime})$ is drawn according to $\mathbb{D}^{2m}$. Therefore, the expected value in (3.2) does not change if the random sample $z_{i}$ and the random sample $z^{\prime}_{i}$ are mutually exchanged. Formally, this means that the term $(\ell(h,z^{\prime}_{i})-\ell(h,z_{i}))$ is replaced by $-(\ell(h,z^{\prime}_{i})-\ell(h,z_{i}))$. Hence, for any $\underline{\sigma}=(\sigma_{1},\dots,\sigma_{m})\in\{\pm 1\}^{m}$ the expected value in (3.2) is equal to
@@
Since this holds for every $\underline{\sigma}\in\{\pm 1\}^{m}$, it also holds if the components of $\underline{\sigma}$ are distributed uniformly at random according to the uniform distribution over {±$\{\pm 1\}$, denoted $\mathcal{U}_{\pm}$. Hence, the expected value in (3.2) is also equal to
@@
and it can be written as an average of independent random variables, each of which takes values in $[-1,1]$ (see Lemma A.1). Thus, we can apply Hoeffding’s Inequality (cf. Pollard [25, Appendix B, Corollary 3]) to obtain
@@
$\mathbb{D}^{m}\left(\left\{\underline{z}\in\mathcal{Z}^{m}\left|\ \sup\limits_{h\in\mathcal{H}}\big{|}\operatorname{er}_{\mathbb{D}}(h)-\hat{\operatorname{er}}_{\underline{z}}(h)\big{|}\leq\varepsilon_{0}(m,\delta)\right.\right\}\right)\geq 1-\delta,$
@@
The proof of Theorem 3.6 relies on bounding the expected value of $U$ by the expected value of $V$ (see (3.1), (3.2) and (3.3)). Thus, assuming $\mathcal{H}$ to be well-behaved is essential for Theorem 3.6, whereas it is not directly apparent that this assumption is necessary for establishing Theorem 3.8. However, we apply Theorem 3.6 in the proof of Theorem 3.8, which in turn yields the implication (1$\Rightarrow$(2). Hence, the assumption on $\mathcal{H}$ to be well-behaved in the Fundamental Theorem of Statistical Learning, arises from Theorem 3.6.
The arguments for proving the following theorem are inspired by the steps in [27, page 51].
@@
$\mathbb{D}^{m}\left(\left\{\underline{z}\in\mathcal{Z}^{m}\left|\ \sup\limits_{h\in\mathcal{H}}\big{|}\operatorname{er}_{\mathbb{D}}(h)-\hat{\operatorname{er}}_{\underline{z}}(h)\big{|}\leq\varepsilon\right.\right\}\right)\geq 1-\delta.$
Note that this probability is well-defined, as $m\geq m_{0}\geq m_{\mathcal{H}}$ implies the measurability of the map $U(\mathcal{H},m,\mathbb{D})$ from Definition 2.4. Since $m\geq m_{0}^{(3)}$, applying [27, Lemma A.2] yields
@@
As $12m\geq 2m_{0}^{(1)}=d+$ 1, we can apply Lemma 2.6 to obtain
@@
$\left\{\underline{z}\in\mathcal{Z}^{m}\left|\ \sup\limits_{h\in\mathcal{H}}\big{|}\operatorname{er}_{\mathbb{D}}(h)-\hat{\operatorname{er}}_{\underline{z}}(h)\big{|}\leq\varepsilon_{0}(m,\delta)\right.\right\}$
@@
$\mathbb{D}^{m}\left(\left\{\underline{z}\in\mathcal{Z}^{m}\left|\ \sup\limits_{h\in\mathcal{H}}\big{|}\operatorname{er}_{\mathbb{D}}(h)-\hat{\operatorname{er}}_{\underline{z}}(h)\big{|}\leq\varepsilon\right.\right\}\right)$
@@
$\mathbb{D}^{m}\left(\left\{\underline{z}\in\mathcal{Z}^{m}\left|\ \sup\limits_{h\in\mathcal{H}}\big{|}\operatorname{er}_{\mathbb{D}}(h)-\hat{\operatorname{er}}_{\underline{z}}(h)\big{|}\leq\frac{\varepsilon}{4}\right.\right\}\right)\geq 1-\delta$
and $\hat{\operatorname{er}}_{\underline{z}}(\mathcal{A}(\underline{z}))-\hat{\operatorname{opt}}_{\underline{z}}(\mathcal{H})\leq\tfrac{\varepsilon}{4}$. Note that the probability is well-defined, as $m\geq m_{0}\geq m_{\mathcal{H}}$ implies the measurability of $U(\mathcal{H},m,\mathbb{D})$. Fix $m\geq m_{0}$ and $\mathbb{D}\in\mathcal{D}$. By the definition of $\operatorname{opt}_{\mathbb{D}}(\mathcal{H})$, we can choose a hypothesis $h_{\varepsilon}\in\mathcal{H}$ such that $\operatorname{er}_{\mathbb{D}}(h_{\varepsilon})-\operatorname{opt}_{\mathbb{D}}(\mathcal{H})\leq\tfrac{\varepsilon}{4}.$ Now let $\underline{z}\in\mathcal{Z}^{m}$ such that $\sup\limits_{h\in\mathcal{H}}\big{|}\operatorname{er}_{\mathbb{D}}(h)-\hat{\operatorname{er}}_{\underline{z}}(h)\big{|}\leq\frac{\varepsilon}{4}$ to obtain
$\forall h\in\mathcal{H}\colon\;\big{|}\operatorname{er}_{\mathbb{D}}(h)-\hat{\operatorname{er}}_{\underline{z}}(h)\big{|}\leq\frac{\varepsilon}{4}.$ (3.6)
@@
$\leq\hat{\operatorname{er}}_{\underline{z}}(h_{\varepsilon})+\frac{\varepsilon}{2}$ (definition of $(\text{definition of $\hat{\operatorname{opt}}_{\underline{z}}(\mathcal{H})$})$
@@
$\leq\operatorname{opt}_{\mathbb{D}}(\mathcal{H})+\varepsilon$ (choice of $(\text{choice of $h_{\varepsilon}$}).$
@@
$C:=\;$$\left\{\underline{z}\in\mathcal{Z}^{m}\left|\ \sup\limits_{h\in\mathcal{H}}\big{|}\operatorname{er}_{\mathbb{D}}(h)-\hat{\operatorname{er}}_{\underline{z}}(h)\big{|}\leq\frac{\varepsilon}{4}\right.\right\}$
@@
(4$\Rightarrow$(1)
To complete the proof of the Fundamental Theorem of Statistical Learning, we now justify that hypothesis spaces with infinite VC dimension are not PAC learnable. Our arguments are guided by the proof of the No Free Lunch Theorem presented in [27, Theorem 5.1].
@@
is bounded and $\mathcal{P}(\mathcal{Z}_{S}^{m})$–measurable for any $i\in[T]$. Next, we justify inequality (3.9). There are $k:=|S^{m}|=(2m)^{m}$ possible $m$–tuples of instances from $S$. We write $S^{m}=\{\underline{a}_{1},\dots,\underline{a}_{k}\}$ with $\underline{a}_{j}\in S^{m}$. Further, given $j\in[k]$, $\underline{a}_{j}=(x_{1},\dots,x_{m})$ and $i\in[T]$, we write
@@
=$=\begin{cases}(\frac{1}{2m})^{m}=\frac{1}{k}&\text{if }\underline{z}=\underline{z}_{j}^{i}\text{ for some }j\in[k],\\ 0&\text{otherwise}.\end{cases}$
@@
[…]
@@
=$=\begin{cases}(\frac{1}{2m})^{m}=\frac{1}{k}&\text{if }\underline{z}=\underline{z}_{j}^{i}\text{ for some }j\in[k],\\ 0&\text{otherwise}.\end{cases}$
@@
Thus, applying Markov’s Inequality ([27, Appendix B.1]) yields
@@
=\,= $\hat{\mathbb{D}}^{m}(\{\underline{z}\in\mathcal{Z}^{m}\mid\operatorname{er}_{\mathbb{D}}(\mathcal{A}(\underline{z}))>\tfrac{1}{8}\})$
[…]

### v3

@@ 3 Fundamental Theorem of Statistical Learning
The aim of this section is to establish the Fundamental Theorem of Statistical Learning. This result, due to Blumer, Ehrenfeucht, Haussler and Warmuth [5, Theorem 2.1 (i)], beautifully relates the notions VC dimension and PAC learnability. More precisely, it states that, under certain broadly applicable measurability conditions, a hypothesis space is PAC learnable if and only if its VC dimension is finite. In the proofs presented in this article, we follow modern expositions such as Anthony and Bartlett [2], Shalev-Schwartz and Ben-David [28] and Vidyasagar [36]. While the technical steps of our arguments are largely the same as (and in parts identical to) those in these references, our emphasis is on identifying the measurability conditions required for the results to remain valid. We present a self-contained and detailed proof in a very general set-up, explicitly indicating which steps demand measurability assumptions. We thus extract the minimal measurability requirements, deliberately avoiding unnecessarily strong assumptions.
@@
$\underline{z}$ $\mapsto U(\underline{z}):=\sup\limits_{h\in\mathcal{H}}\big|\hskip-1.25pt\operatorname{er}_{\mathbb{D}}(h)-\hat{\operatorname{er}}_{\underline{z}}(h)\big|$
@@
The proof of the Fundamental Theorem of Statistical Learning presented below, specifically the proof of Theorem 3.6, relies on computing and bounding the expected values of the maps $U$ and $V$. Therefore, the measurability of these maps is necessary for our arguments to apply. In other words, the notion of well-behavedness reflects the minimal measurability requirements of the Fundamental Theorem of Statistical Learning, and thus gives rise to maximally general results. In particular, well-behavedness avoids constraints such as compactness and completeness conditions. Sufficient conditions – such as permissibility and universal separability – for hypothesis spaces to be well-behaved are discussed in Appendix A.2. In fact, most hypothesis spaces arising in applications satisfy these conditions.
@@
At this point, we emphasize that in the literature there are several extended versions of this theorem. These entail further equivalences and quantitative aspects such as sample complexity bounds (see e.g. [2, Theorem 5.5], [5, Theorem 2.1], [28, Theorem 6.7 and Theorem 6.8]). The main aim for the remainder of this section is to establish the Fundamental Theorem of Statistical Learning as stated above by virtue of the following more extended version:
@@
[…]
@@
Let $\mathcal{X}$ be a non-empty set, let $\Sigma_{\mathcal{Z}}$ be a $\sigma$–algebra on $\mathcal{Z}=\mathcal{X}\times\{0,1\}$ and let $h\in\{0,1\}^{\mathcal{X}}$ with $\Gamma(h)\in\Sigma_{\mathcal{Z}}$. Then the hypothesis space $\mathcal{H}=\{h\}$ has the uniform convergence property with respect to any set $\mathcal{D}$ of distributions on $(\mathcal{Z},\Sigma_{\mathcal{Z}})$.
@@
$\mathbb{D}^{m}\left(\left\{\underline{z}\in\mathcal{Z}^{m}\left|\ \big|\hskip-1.25pt\operatorname{er}_{\mathbb{D}}(h)-\hat{\operatorname{er}}_{\underline{z}}(h)\big|>\varepsilon\right.\right\}\right)\leq\delta.$
@@
$\big|\hskip-1.25pt\operatorname{er}_{\mathbb{D}}(h)-\hat{\operatorname{er}}_{\underline{z}}(h)\big|=\frac{1}{m}\big|Y_{1}(\underline{z})+\dots+Y_{m}(\underline{z})\big|,$
@@
is a random variable [^9] The map $Y_{i}$ is a translation of the map $\theta_{i}\colon\mathcal{Z}^{m}\to\{0,1\},\ \underline{z}\mapsto\ell(h,z_{i})$, which is $\Sigma_{\mathcal{Z}}^{m}$–measurable, as $\Gamma(h)\in\Sigma_{\mathcal{Z}}$ implies $\theta_{i}^{-1}(\{0\})=\mathcal{Z}\times\dots\times\mathcal{Z}\times\underbrace{\Gamma(h)}_{i\text{--th position}}\times\mathcal{Z}\times\dots\times\mathcal{Z}\in\Sigma_{\mathcal{Z}}^{m}.$ with $\mathbb{E}_{\underline{z}\sim\mathbb{D}^{m}}[Y_{i}(\underline{z})]=0$ for any $i\in[m]$. Straightforward computations show that the random variables $Y_{1},\dots,Y_{m}$ are independent. Thus, we can apply Hoeffding’s Inequality (cf. Pollard [26, Appendix B, Corollary 3]) to obtain
[…]
@@
We next prove a result that shows how the growth function $\pi_{\mathcal{H}}$ can be used to bound the probability involved in the definition of the uniform convergence property. The proof requires the hypothesis space to be well-behaved, since its core lies in bounding the expected values of the random variables $U$ and $V$. The original proof of Vapnik and Chervonenkis is also based on the assumption that these maps are measurable (see [35, pages 265 and 268]). We present here a modified and adapted version of [28, proof of Theorem 6.11], additionally highlighting explicitly where the required measurability assumptions come into play. While in the literature similar and even sharper bounds are established under the stronger assumption of a permissible hypothesis space (cf. Haussler [15, Theorem 2 and Corollary 1], Pollard [26, Chapter II]), our result applies to the more general framework of well-behaved hypothesis spaces, thereby extending the scope of the results beyond the permissible setting.
@@
$\mathbb{D}^{m}\left(\left\{\underline{z}\in\mathcal{Z}^{m}\left|\ \sup\limits_{h\in\mathcal{H}}\big|\hskip-1.25pt\operatorname{er}_{\mathbb{D}}(h)-\hat{\operatorname{er}}_{\underline{z}}(h)\big|\leq\varepsilon_{0}(m,\delta)\right.\right\}\right)\geq 1-\delta,$
@@
$\underline{z}$ $\mapsto U(\underline{z}):=\sup\limits_{h\in\mathcal{H}}\big|\hskip-1.25pt\operatorname{er}_{\mathbb{D}}(h)-\hat{\operatorname{er}}_{\underline{z}}(h)\big|.$
We have to show that $\mathbb{D}^{m}(\{\underline{z}\in\mathcal{Z}^{m}\mid U(\underline{z})\leq\varepsilon_{0}(m,\delta)\})\geq 1-\delta$. Since the random variable $U$ is non-negative, by Markov’s Inequality (see [28, Appendix B.1]) it suffices to establish that
@@
writing $\underline{z}=(z_{1},\dots,z_{m})$ and $\underline{z}^{\prime}=(z^{\prime}_{1},\dots,z^{\prime}_{m})$. The expected value in (3.2) is determined based on a choice of multi-sample $(\underline{z},\underline{z}^{\prime})\in\mathcal{Z}^{2m}$ whose components are independently distributed according to $\mathbb{D}$, i.e. $(\underline{z},\underline{z}^{\prime})$ is drawn according to $\mathbb{D}^{2m}$. Therefore, the expected value in (3.2) does not change if the random sample $z_{i}$ and the random sample $z^{\prime}_{i}$ are mutually exchanged. Formally, this means that the term $(\ell(h,z^{\prime}_{i})-\ell(h,z_{i}))$ is replaced by $(\ell(h,z^{\prime}_{i})-\ell(h,z_{i}))$. Hence, for any $\underline{\sigma}=(\sigma_{1},\dots,\sigma_{m})\in\{\pm 1\}^{m}$ the expected value in (3.2) is equal to
@@
Since this holds for every $\underline{\sigma}\in\{\pm 1\}^{m}$, it also holds if the components of $\underline{\sigma}$ are distributed uniformly at random according to the uniform distribution over $\{\pm 1\}$, denoted $\mathcal{U}_{\pm}$. Hence, the expected value in (3.2) is also equal to
@@
and it can be written as an average of independent random variables, each of which takes values in [1,1][-1,1] (see Lemma A.1). Thus, we can apply Hoeffding’s Inequality (cf. Pollard [26, Appendix B, Corollary 3]) to obtain
@@
$\mathbb{D}^{m}\left(\left\{\underline{z}\in\mathcal{Z}^{m}\left|\ \sup\limits_{h\in\mathcal{H}}\big|\hskip-1.25pt\operatorname{er}_{\mathbb{D}}(h)-\hat{\operatorname{er}}_{\underline{z}}(h)\big|\leq\varepsilon_{0}(m,\delta)\right.\right\}\right)\geq 1-\delta,$
@@
[…]
The arguments for proving the following theorem are inspired by the steps in [28, page 51].
@@
$\mathbb{D}^{m}\left(\left\{\underline{z}\in\mathcal{Z}^{m}\left|\ \sup\limits_{h\in\mathcal{H}}\big|\hskip-1.25pt\operatorname{er}_{\mathbb{D}}(h)-\hat{\operatorname{er}}_{\underline{z}}(h)\big|\leq\varepsilon\right.\right\}\right)\geq 1-\delta.$
Note that this probability is well-defined, as $m\geq m_{0}\geq m_{\mathcal{H}}$ implies the measurability of the map $U(\mathcal{H},m,\mathbb{D})$ from Definition 2.4. Since $m\geq m_{0}^{(3)}$, applying [28, Lemma A.2] yields
@@
As $2m\geq 2m_{0}^{(1)}=d+1$, we can apply Lemma 2.6 to obtain
@@
[…]
@@
$\mathbb{D}^{m}\left(\left\{\underline{z}\in\mathcal{Z}^{m}\left|\ \sup\limits_{h\in\mathcal{H}}\big|\hskip-1.25pt\operatorname{er}_{\mathbb{D}}(h)-\hat{\operatorname{er}}_{\underline{z}}(h)\big|\leq\frac{\varepsilon}{4}\right.\right\}\right)\geq 1-\delta$
and $\hat{\operatorname{er}}_{\underline{z}}(\mathcal{A}(\underline{z}))-\hat{\operatorname{opt}}_{\underline{z}}(\mathcal{H})\leq\tfrac{\varepsilon}{4}$. Note that the probability is well-defined, as $m\geq m_{0}\geq m_{\mathcal{H}}$ implies the measurability of $U(\mathcal{H},m,\mathbb{D})$. Fix $m\geq m_{0}$ and $\mathbb{D}\in\mathcal{D}$. By the definition of $\operatorname{opt}_{\mathbb{D}}(\mathcal{H})$, we can choose a hypothesis $h_{\varepsilon}\in\mathcal{H}$ such that $\operatorname{er}_{\mathbb{D}}(h_{\varepsilon})-\operatorname{opt}_{\mathbb{D}}(\mathcal{H})\leq\tfrac{\varepsilon}{4}.$ Now let $\underline{z}\in\mathcal{Z}^{m}$ such that $\sup\limits_{h\in\mathcal{H}}\big|\hskip-1.25pt\operatorname{er}_{\mathbb{D}}(h)-\hat{\operatorname{er}}_{\underline{z}}(h)\big|\leq\frac{\varepsilon}{4}$ to obtain
$\forall h\in\mathcal{H}\colon\;\big|\hskip-1.25pt\operatorname{er}_{\mathbb{D}}(h)-\hat{\operatorname{er}}_{\underline{z}}(h)\big|\leq\frac{\varepsilon}{4}.$ (3.6)
@@
$\leq\hat{\operatorname{opt}}_{\underline{z}}(\mathcal{H})+\frac{\varepsilon}{2}$ $(\text{$\mathcal{A}$ is NMSE})$
@@
$\leq\operatorname{opt}_{\mathbb{D}}(\mathcal{H})+\varepsilon$ $(\text{choice of $h_{\varepsilon}$}).$
@@
$\subseteq\;$ $\left\{\underline{z}\in\mathcal{Z}^{m}\mid\operatorname{er}_{\mathbb{D}}(\mathcal{A}(\underline{z}))-\operatorname{opt}_{\mathbb{D}}(\mathcal{H})\leq\varepsilon\right\}\!.$
@@
[…]
To complete the proof of the Fundamental Theorem of Statistical Learning, we now justify that hypothesis spaces with infinite VC dimension are not PAC learnable. Our arguments are guided by the proof of the No Free Lunch Theorem presented in [28, Theorem 5.1].
@@
is bounded and $\mathcal{P}(\mathcal{Z}_{S}^{m})$–measurable for any $i\in[T]$. Next, we justify inequality (3.9). There are $k:=|S^{m}|=(2m)^{m}$ possible mm–tuples of instances from $S$. We write $S^{m}=\{\underline{a}_{1},\dots,\underline{a}_{k}\}$ with $\underline{a}_{j}\in S^{m}$. Further, given $j\in[k]$, $\underline{a}_{j}=(x_{1},\dots,x_{m})$ and $i\in[T]$, we write
@@
$=\begin{cases}(\frac{1}{2m})^{m}=\frac{1}{k}&\text{if }\underline{z}=\underline{z}_{j}^{i}\text{ for some }j\in[k],\\ 0&\text{otherwise}.\end{cases}$
@@
$\geq\;$ $\min\limits_{j\in[k]}\underbrace{\frac{1}{T}\sum\limits_{i=1}^{T}\operatorname{er}_{\mathbb{D}_{i}}(\mathcal{A}_{S}(\underline{z}_{j}^{i}))}_{\geq\frac{1}{2}\underbrace{\min\limits_{r\in[p]}\frac{1}{T}\underbrace{\sum\limits_{i=1}^{T}\mathds{1}_{\mathcal{Z}_{S}\setminus\Gamma(f_{i})}({v_{r}},(\mathcal{A}_{S}(\underline{z}_{j}^{i}))({v_{r}}))}_{=\frac{T}{2}}}_{=\frac{1}{2}}}$
$\geq\;$ $\frac{1}{4}.\qed$
@@
$=\begin{cases}(\frac{1}{2m})^{m}=\frac{1}{k}&\text{if }\underline{z}=\underline{z}_{j}^{i}\text{ for some }j\in[k],\\ 0&\text{otherwise}.\end{cases}$
@@
Thus, applying Markov’s Inequality ([28, Appendix B.1]) yields
@@
$=\,$ $\hat{\mathbb{D}}^{m}(\{\underline{z}\in\mathcal{Z}^{m}\mid\operatorname{er}_{\mathbb{D}}(\mathcal{A}(\underline{z}))>\tfrac{1}{8}\})$
$\geq\,$ $\frac{\mathbb{E}_{\underline{z}\sim\hat{\mathbb{D}}^{m}}[\operatorname{er}_{\mathbb{D}}(\mathcal{A}(\underline{z}))]-\tfrac{1}{8}}{1-\tfrac{1}{8}}$
$\geq\,$ $\frac{\tfrac{1}{4}-\tfrac{1}{8}}{\tfrac{7}{8}}=\frac{1}{7}.\qed$